OpenStudy (anonymous):

If r = 9, b = 5, and g = -6, what does (r + b - g)(b + g) equal? -20 -8 8 20

OpenStudy (mathmale):

You might find it just slightly easier to evaluate (r + b - g) and (b + g) separately, and then to multiply your two results together. b+g = 5-6=??